OpenStudy (anonymous):

write the decimal as a fraction or a mixed number in simplest form 3.4

OpenStudy (anonymous):

3 and 4/10 which is 2/5.. 3 2/5 =17/5

OpenStudy (anonymous):

17/5 is not right tho, lots of people are answering it as 3 2/5 which is also wrong , the right answer would be 3 4/9 , why tho?

OpenStudy (anonymous):

theanswer is3and 2/5 which can also be written asan improperfraction 17/5. (so itis thesame answer).

OpenStudy (anonymous):

31/9 is 3.4 so i guess thats how you get 3 4/9

OpenStudy (anonymous):

see check this out!

OpenStudy (anonymous):

it is not simply a 4. it has a bar on top, representing that it repeats the 4. so it is really 3.444444: simplified as 3 and 4/9 because 44(just adding on anohter 4 to make it in the hundred column) / 100 ( 44/100= 4/9